OpenStudy (konradzuse):

evaluate each integral (lnx)^3 / (x) dx

OpenStudy (anonymous):

since the d/dx of ln (x) is the the denominator the integral should look like this \[\int\limits_{}^{}(\ln(x))^3d(\ln(x))=\int\limits_{}^{}(\ln(x))^3/xdx\] so because d/dx(ln(x)) is sitting right there you can think of it as "u" so (ln(x))^3=u^3 INT u^3=u^4/4 since u=ln(x) (ln(x))^4/4+C